OpenStudy (anonymous):

Solve the following problems completely . a. Christian can do a job in 6 hours. Requesting Boyet to work with him, the two can do the jon in 4 1/2 hours. How long would it take Boyet to do the job alone? b. A motorcycle can maintain constant speed of 16 kilometers per hour relative to the water. The boat makes a trip upstream to a certain point is 20 minutes; the return trip take 15 minutes, What is the speed of the current? c. The perimeter of a rectangle is 28 cm. IF the length were subtracted by2 and the width added by 2, the rectangle would have been a square. What are the dimens

OpenStudy (anonymous):

a. 6 + 4 1/2 ÷ 2 = ? urg I did the other 2 questions earlier this year D: and I forgot how to do them xD

OpenStudy (anonymous):

Nooooo D; Upou?

OpenStudy (anonymous):

Upou? Idk what that means

OpenStudy (anonymous):

Nevermind. T_T but please do remember :<

OpenStudy (anonymous):

Give me a sec and I'll see if I can find them in my coursework. Then I shall attempt to help you.

OpenStudy (anonymous):

Thanks :)

OpenStudy (anonymous):

Okay it wasn't the exact same problem Here's what I had and how I worked it out A boat travels 30 mph. A patrol boat starts 4 hours later from the same place and travels at 35 mph in the same direction. How long will the patrol boat require to overtake the first boat? Step 1: Let t = time to overtake by patrol boat. t + 4 = time the first boat is traveling. rate time distance first boat: 30 t + 4 30(t + 4) patrol boat: 35 t 35t At the time of overtake, the distances of the two boats are equal. Step 2: 30(t + 4) = 35t Step 3: Solve 30(t + 4) = 35t. 30t + 120 = 35t 120 = 5t t = 24 hours lol idk if that will help but it was worth a shot.

OpenStudy (anonymous):

Thanks, but I need speed of the current. Man this is hard ! :/

OpenStudy (anonymous):

I'll go post your question in chat :D

OpenStudy (zepp):

a. Christian can do a job in 6 hours. Requesting Boyet to work with him, the two can do the jon in 4 1/2 hours. How long would it take Boyet to do the job alone? Okay, let the Christian's speed by x, and Boyet's speed y. Christian can do the job in 6 hours, so the whole work can be expressed as \(6x\); If they do it together, it will take 4 1/2 hours, whole work can be expressed as \(4\frac{1}{2}(x+y)\) You now a system, solve for x :)

OpenStudy (zepp):

Why is the whole work equal to the speed times the time required to do it? Let's see \(\large \frac{\text{Whole work}}{\text{Speed}}=\text{Time}\) Let's play around a little bit with that \(\large \text{Whole work}=\text{Time} *\text{Speed}\)

OpenStudy (zepp):

So let's solve it; Since the whole work is equal, then the two equations above are equal \(\large \begin{align} 6x&=4\frac{1}{2}(x+y) \\ 6x\div 4\frac{1}{2}&=(x+y) \\ 3x&=x+y \\ 3x-x&=y \\ 2x&=y \end{align}\)

OpenStudy (zepp):

Good, \(2x=y\), what does it mean?

OpenStudy (zepp):

I don't know either. :P

OpenStudy (zepp):

Haha, that mean that y, Boyet's speed, is twice x, Christian's speed.

OpenStudy (zepp):

Therefore, Boyet would take half of Christian's time to do the job, 3 hours.

OpenStudy (zepp):

Is that clear enough? We can go to the second question?

OpenStudy (anonymous):

Yep loud and clear.

OpenStudy (zepp):

b. A motorcycle can maintain constant speed of 16 kilometers per hour relative to the water. The boat makes a trip upstream to a certain point is 20 minutes; the return trip take 15 minutes, What is the speed of the current? I don't quite understand this one, a motorcycle goes in the water? o.o